Match the expression with the correct answer. You may use each answer more than once.
A. ln e
-b. ln 1
c. log 10
d. e^0


A. 0
B. 1
C. none of the above



Answer by mathdoc314(58)

I don't know how you can have this question in front of you and not the information you need to answer it!
Here is some relevant information:
Facts:
1. Logarithm base e is written as ln.
2. Logarithm base 10 is written as log.
3. Any number raised to the power of 0 is 1.
4. Any number raised to the power of 1 is itself.
5. The logarithm function is the inverse of the exponential function.
That means 10^(log x) = x and e^(ln x) = x for x>0,
and also
log(10^x) = x and ln(e^x) = x for any x.
More facts: e^0 = 1, e^1 = e, 10^1 = 10, 10^0 = 0
e^(ln 1) = 1
e^(ln e) = e
10^(ln 10) = 1
10^(ln 1) = 0
If you look closely, the answer is here.